Operations Report – February 17th

Crew 275 Operations Report 17Feb2023
SOL: 5

Name of person filing report: Quentin ROYER
Non-nominal systems:
• Toilet
• Science Dome: 3 leaks in the walls – same as yesterday
• Propane Generator

Notes on non-nominal systems:
• Propane generator: power outage from [unknown] to 5:32 am

Spirit rover used: No
• Hours: 209.0
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Opportunity rover used: No
• Hours: 113.0
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Curiosity rover used: No
• Hours: 220.9
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Perseverance rover used: No
• Hours: 255.4
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

General notes on rovers: Today’s EVA was postponed for safety reason after the propane alarm in the morning, because the crewmembers were tired, and the Crew wanted to stay at the Hab to monitor any unexpected situation during the day. Indeed, crewmembers planned for this EVA included the Commander and the Crew Engineer.

Summary of Hab operations: Power outage in the morning (see below the section “Summary of health and safety issues”). The weather station which will be deployed during the next EVA was fully assembled and tested in the Lower Deck.
• WATER USE: 14.42 gallons
• Water: 396.54 gallons
• Static tank pipe heater: On
• Static tank heater: On
• Toilet tank emptied: No

Summary of internet: Hughes Net was used during the day before the opening of COMMS for email purposes, and Starlink (MDRSCampus-Guest) is used during COMMS.

Summary of suits and radios: Nothing to report as today’s EVA was postponed to another day.

Summary of GreenHab operations: The Crew botanist conducted routine operations.
• WATER USE: 7.39 gallons
• Heater: On
• Supplemental light: From 10:00 pm to 2:00 am
• Harvest: 4 g of basil – 3 g of thyme

Summary of ScienceDome operations: The Commander and Crew Scientist tested some experimental protocols.
• Dual split: Heat / From 10 pm to 7 am

Summary of RAM operations: The RAM was not used because of the freezing temperatures. There is no more whistling sound. There is no more leak.

Summary of any observatory issues: The robotic observatory was open until around 12:30.

Summary of health and safety issues:

At 5:27 am, the Commander woke up and detected that there was a power outage, and a very low temperature on the Upper and Lower Decks (56°F / 13°C on the Upper Deck, much less on the Lower Deck). He immediately notified Mission Support who promptly reacted. He also notified the Crewmembers to wear hot clothes and to stay in their rooms, where the temperature was warmer.

At 5:32 am, power came back in the main Hab, and propane detectors warmed up.

At 5:34 am, the Lower Deck propane detector’s alarm rang, announcing a “possible gas leak”. The Commander, the HSO and the Crew Engineer reacted to the emergency. The Commander notified Mission Support. The HSO and Crew Engineer went to the RAM to get the two gas detectors.

Back in the Hab, crewmembers began an inspection of the Lower Deck with the gas detectors.

At 5:39 am, the alarm stopped. Crew members continued to monitor the situation, and seek for a potential gas leak. As the temperature was very low, Mission Support instructed to Crew to switch on the Lower Deck heater. This took about 20 minutes, because the heater did not switch on before (there was no gas going out from the pipe, even though the valve was open).

We finally managed to switch on the heater, and Crew members rested before starting their daily activities at 6:45. A propane gas detector was left ON in the Lower Deck to monitor the situation. It ran out of battery around 2 hours later.

The yellow gas detector showed weird values: when the red gas detector was beeping, the yellow one still showed the same value of LEL (0 to 2%)
The heater was left on during most of the day, and the temperature was back to normal (72°F) at 10:57 am in the Upper Deck.

Questions, concerns, and requests to Mission Support:
• The oven is working now. Same for the heater on the lower deck, after 20 minutes of trying to switch it on, it switched on
• Greenhab Radio on channel 1: the sound of the radio is very low, even though the volume is at MAX, it is hard to hear the person speaking, what can we do?
• Both gas detectors ran out of battery, each one work with 3 AAA batteries, and we have no more AAA batteries.

Copyright © The Mars Society. All rights reserved. | Main Site